OpenStudy (dreamie98):

What is the prime factor of x^6 + 1 ?

OpenStudy (hitch):

\[x ^{6}+1 = (x ^{2}+x+1)(x ^{2}−x+1)(x+1)(x−1)\]

zepdrix (zepdrix):

Hey Dreamie :) We can make use of our `Sum of Cubes Identity`: \(\large\rm a^3+b^3=(a+b)(a^2-ab+b^2)\) We need to rewrite this 6th power as a cube though. We can do this little trick using rules of exponents,\[\large\rm x^6=(x^2)^3\]Ahh that should help! So our expression,\[\large\rm x^6+1\]is really,\[\large\rm (x^2)^3+1^3\]and we can apply our identity, yay!\[\large\rm =(x^2+1)\left((x^2)^2-(x^2)(1)+1^2\right)\]

zepdrix (zepdrix):

So we have,\[\large\rm (x^2+1)(x^4-x^2+1)\]Which uhhh, ya I think both of those are prime factors. I'm not quite sure how Hitch factored it down so far >.< hmm

OpenStudy (hitch):

it is true by the end Mr.Zepdrix

OpenStudy (dreamie98):

thanks for the help guys, I couldn't answer because my internet went out D: <3